# Universal bound on the Lyapunov spectrum of quantum master equations

### v1

@@ Abstract
The spectral properties of positive maps provide pivotal information for understanding the dynamics of quantum systems interacting with their environment. Furthermore, central problems in quantum information such as the characterization of entanglement may be reformulated in terms of spectral properties of positive maps. The present work aims to contribute to a better understanding of the spectrum of positive maps. Specifically, our main result is a new proof of a universal bound on the $d^{2}-1$ generically non vanishing decay rates $\Gamma_{i}$ of time-autonomous quantum master equations on a $d$-dimensional Hilbert space:
$\Gamma_{\mathrm{max}}\,\leq\,c_{d}\,\sum_{i=1}^{d^{2}-1}\Gamma_{i}$
The prefactor $c_{d}$ , which we explicitly determine, depends only on the dimension $d$ and varies depending on the sub-class of positive maps to which the semigroup solution of the master equation belongs. We provide a brief but self-consistent survey of these concepts. We obtain our main result by resorting to the theory of Lyapunov exponents, a central concept in the study of dynamical systems, control theory, and out-of-equilibrium statistical mechanics. We thus show that progress in understanding positive maps in quantum mechanics may require ideas at the crossroads between different disciplines. For this reason, we adopt a notation and presentation style aimed at reaching readers with diverse backgrounds.
@@ 1 Introduction
Physics is an experimental science. A recurring question is how to extract dynamical laws from experimental data. In the case of open quantum systems, this question is often formulated by asking whether an observed evolution (dynamical map) can be mathematically understood as a family of linear maps that enjoys the semigroup property and whose members are all positivity preserving. In that case, there also exists a precise and computable answer for time-autonomous dynamics [WoEiCuCi2008]. Specifically, one needs to check whether the logarithm of a tomographically estimated dynamical map at a given time, i.e., a single snapshot of the dynamical map, satisfies the necessary and sufficient conditions characterizing generators of completely positive semigroups and, therefore, solutions of the Lindblad [LinG1976] Gorini-Kossakowski-Sudarshan [GoKoSu1976] master equation. These conditions are preservation of the trace and self-adjointness, as well as conditional complete positivity [EvaD1977] (see also [WolM2012, § 7]) of the generator. Whilst the answer is mathematically complete and constructive, actual implementations encounter several stumbling blocks. To start with, the complex logarithm is a multi-valued function so that, in general, checks may require to extend the search to multiple branches. In addition, the search must be robust against state preparation and measurement errors that occur in realistic situations. Finally, the computational cost grows exponentially with the number of dimensions. Connecting to this cost, it is also worth mentioning that it is in fact possible to prove [CuEiWo2012] that the 11-IN-33-SAT, an instance of a NP-complete problem in computational complexity theory [MoMe2011, § 5], is amenable by a Turing reduction to proving that the exponential of a certain non-representational (i.e., abstract and non related to any physical model [IpCa2021]) generator yields an element of a complete positive semigroup. This means that in principle assessing whether the logarithm of a dynamical map specifies the generator of a completely positive semigroup is at least as hard as solving any NP problem on a classical computer. In practice and for systems with few degrees of freedom, the result of [WoEiCuCi2008] is the basis for computationally viable algorithms which are relevant for to-date quantum computing hardware [LiSeKoOnCu2025].
@@
The difficulties discussed above motivate the interest for partial but practically more direct answers that provide necessary conditions (“witnesses”) to identify a completely positive semigroup. It was conjectured long ago [KimG2002] that complete positivity of a homogeneous semigroup of dynamical maps necessarily entails an upper bound on the relaxation rates characterizing the contractive forward evolution. Relaxation rates are commonly measured in laboratory experiments. The conjecture was later extensively corroborated by proofs under additional hypotheses and examples exhibiting tightness of the putative bound [ChKiKoSh2021]. Mathematically, relaxation rates are the real parts of the logarithms of the eigenvalues of the flow solution of a linear differential equation. If the differential equation is time-autonomous, the forward in time flow it generates constitutes a homogeneous semigroup and the relaxation rates coincide with the Lyapunov exponents [AdrL1995]. These considerations led to a first proof of the conjecture by combining ideas of dynamical systems and open quantum system theory [MGKiCh2025]. The proof, however, left open the question of tight bounds for the relaxation rates of positive but non completely positive maps. This is an important question because positive (i.e., positivity-preserving) linear maps—which are fundamental to quantum mechanics—have a far more articulated structure than those appearing in the theory of classical stochastic processes [ChKo2009]. In [ChvEnKiMG2025, vEnChKiMG2025] we were able to provide a complete answer to this question using different and more traditional methods of algebraic quantum theory. Specifically, the results of [ChvEnKiMG2025, vEnChKiMG2025] show that the same bound holds for maps that are at least $2$-positive and that this requirement cannot be weakened further. We emphasize that $2$-positive maps are completely positive maps for single-qubit systems. The relaxation rates of simply positive (or $1$-positive) maps only obey the trivial bound dictated by contractivity, as do classical positive maps. Finally, our results in [ChvEnKiMG2025, vEnChKiMG2025] single out the class of Schwarz maps, a subset of positive maps containing all $2$-positive maps, by showing that their relaxation rates obey a weaker yet non trivial bound.
@@
The structure of the paper is as follows. In section 2 we briefly recall basic facts about positive linear maps in classical and quantum physics. In section 3 we recall that any generator of trace- and self-adjointness preserving linear maps embeds a classical master equation [TeMa1992] if and only if the Kossakowski conditions hold [KosA1972, GoKoSu1976] (see also [RiHu2012, § 4.3]). In section 4 we prove that the foliation of the set of positive maps in quantum mechanics in a hierarchy of subsets translates into conditions on the contraction rate of the volume form of state operators. These results are instrumental to prove the main result of this work. Section 5 contains our main result. We include some examples illustrating the bounds in section LABEL:sec:examples. The last section is devoted to conclusions and outlook.
@@ 2.1 Classical master equations
Probability vectors valued on a finite-dimensional configuration space are the main objects in the theory of classical stochastic processes. These are real vectors with all positive entries summing up to one. Geometrically they form a simplex, a convex compact subset of a real Euclidean vector space [BeZy2006, § 2].
If the stochastic process is a differentiable Markov chain on a $d$-dimensional Euclidean configuration space, then the family of probability vectors $\mathsf{p}(t)$ that describes the state of the process as $t$ increases, satisfies a first-order system of differential equations of the form
$\dot{\mathsf{p}}(t)=\mathsf{G}(t)\,\mathsf{p}(t)\,.$ (1)
We refer to the family of $d\,\times\,d$-matrices $\mathsf{G}(t)$ as the (time-dependent) generator of the dynamics. We require $\mathsf{G}(t)$ to depend on the time variable $t$ in a sufficiently regular manner to satisfy the conditions of global existence of solutions [AdrL1995, § 1]. In addition, we require $\mathsf{G}(t)$ to be column-diagonally dominated with positive off-diagonal entries, that is,
$\mathsf{G}_{i,j}(t)\begin{cases}\,\geq\,0&\forall\,i\neq j\\[8.5359pt] =-\sum_{k\neq i}^{d}\mathsf{G}_{k,i}(t)&\forall\,i=j\end{cases}$ (2)
@@
$\mathsf{T}(t,s)=\mathcal{T}\exp\left(\int_{s}^{t}\mathrm{d}u\,\mathsf{G}(u)\right)$ (3)
solving (1) with initial condition
$\mathsf{T}(s,s)=\mathsf{1}_{d}$
@@
$\mathsf{T}(t_{\mathfrak{3}},t_{\mathfrak{1}})=\mathsf{T}(t_{\mathfrak{3}},t_{\mathfrak{2}})\mathsf{T}(t_{\mathfrak{2}},t_{\mathfrak{1}})\qquad\forall\,t_{\mathfrak{3}}\,\geq\,t_{\mathfrak{2}}\,\geq\,t_{\mathfrak{1}}$
@@
$\|\mathsf{T}(t,s)(\mathsf{p}(s)-\tilde{\mathsf{p}}(s))\|_{\mathfrak{1}}\,\leq\,\left\|\mathsf{p}(s)-\tilde{\mathsf{p}}(s)\right\|_{\mathfrak{1}}\,\equiv\,\sum_{i=1}^{d}|\mathsf{p}_{i}(s)-\tilde{\mathsf{p}}_{i}(s)|\,.$ (4)
Combining contractivity with the property of mapping a convex compact set into itself paves the way for the application of Perron-Frobenius theorem (see e.g [SenE2006, § 1]). The upshot is that the spectrum of any of the semigroup elements (3) is confined to the unit disk of the complex plane, and it is symmetric with respect with the real axis. Furthermore, there is at least one eigenvalue equal to one whose right eigenvector is a probability vector after an appropriate choice of normalization.
@@ 2.2 Quantum master equations
State operators are the main objects in the theory of open quantum systems. In the case of a Hilbert space with $d$ dimensions, they are positive semi-definite (henceforth just called “positive”) $d\,\times\,d$ matrices with unit trace. Geometrically they also form a convex compact subspace of the matrix space $\mathcal{M}_{d}(\mathbb{C})$ of $d\,\times\,d$ matrices with complex entries [BeZy2006, § 8.4].
Based only on the requirements of trace and self-adjointness preservation it is in general possible to derive the matrix valued equation
$\dot{\bm{\alpha}}(t)=\mathfrak{L}(t)[\bm{\alpha}(t)]\,.$ (5)
@@
$\begin{split}&\mathfrak{L}(t)[\operatorname{O}]=-\imath\,\left[\operatorname{H}(t),\operatorname{O}\right]+\sum_{\ell=1}^{\mathrm{d}^{2}-1}\mathscr{c}_{\ell}(t)\frac{\left[\operatorname{L}_{\ell}(t)\,,\operatorname{O}\operatorname{L}_{\ell}^{\dagger}(t)\right]+\left[\operatorname{L}_{\ell}(t)\operatorname{O}\,,\operatorname{L}_{\ell}^{\dagger}(t)\right]}{2}\\ &\begin{cases}\operatorname{Tr}\left(\operatorname{L}_{\ell}^{\dagger}(t)\operatorname{L}_{k}(t)\right)=\delta_{\ell,k}\\ \operatorname{Tr}\left(\operatorname{L}_{\ell}(t)\right)=0\\ \operatorname{H}(t)=\operatorname{H}^{\dagger}(t)\end{cases}\hskip 14.22636pt\forall\,t\,\geq\,0\hskip 14.22636pt\&\hskip 14.22636pt\forall\,\ell,k=1,\dots,d^{2}-1\end{split}$ (6)
where the $\mathscr{c}_{\ell}(t)$’s are scalar couplings and the $\operatorname{L}_{\ell}(t)$ are commonly referred to as decoherence or jump operators. As in the classical case, we assume that any explicit time dependence is sufficiently regular to satisfy the conditions for global existence of solutions of (5). In particular, here we require
$|c_{\ell}(t)|\,<\,\infty\qquad\forall\,\ell=1,\dots,d^{2}-1$
We refer to (5), always implying the generator form (6), as “quantum master equation”. The main problem with (5) is that it does not ensure positivity preservation without further assumptions on the scalar couplings $\mathscr{c}_{\ell}(t)$.
A cornerstone result of the theory of open quantum systems is the proof first obtained in the time-autonomous case by Lindblad [LinG1976] and Gorini-Kossakowski-Sudarshan [GoKoSu1976], that if and only if
$\mathscr{c}_{\ell}(t)\,\geq\,0\qquad\forall\,\ell=1,\dots,d^{2}$ (7)
then solutions of (5) can always be couched into the readily positivity-preserving operator sum form (often referred to as Kraus form [KraK1983])
$\Phi(t,s)[\bm{\alpha}(s)]=\sum_{\ell=0}^{d^{2}-1}\operatorname{V}_{\ell}(t,s)\bm{\alpha}(s)\operatorname{V}_{\ell}^{\dagger}(t,s)\qquad\forall\,t\,\geq\,s$ (8)
@@
$\sum_{\ell=0}^{d^{2}-1}\operatorname{V}_{\ell}^{\dagger}(t,s)\operatorname{V}_{\ell}(t,s)=\operatorname{1}_{d}\qquad\forall\,t\,\geq\,s$
@@
It is worth recalling that we can always couch the matrix equation (5) into the form of a vector equation using the vectorization isomorphism relating outer and tensor products [BeZy2006, § 8.2]. Accordingly, a solution of (5) is in one-to-one correspondence with a family of vectors in $\mathbb{C}^{d^{2}}$
$\bm{\alpha}(t)=\sum_{i,j=1}^{d}\mathbf{e}_{i}^{\scriptscriptstyle{(d)}}\mathbf{e}_{j}^{\scriptscriptstyle{(d)}}{}^{\dagger}\alpha_{i,j}(t)\quad\Longleftrightarrow\quad\mathscr{a}(t)=\sum_{i,j=1}^{d}\mathbf{e}_{i}^{\scriptscriptstyle{(d)}}\otimes\mathbf{e}_{j}^{\scriptscriptstyle{(d)}}\alpha_{i,j}(t)$ (9)
where $\left\{\mathbf{e}_{i}^{\scriptscriptstyle{(d)}}\right\}_{i=1}^{d}$ is the canonical (or computational) basis of $\mathbb{C}^{d}$. In words, (9) says that we can construct a vector out of a matrix by stacking rows one above the other. Upon applying this prescription and using the overline symbol $\overline{(\cdot)}\,$ to denote entry-wise complex conjugation, the evolution map (8) takes the familiar form in linear dynamical systems theory (see, e.g., [CeCeVu2009]) of a matrix acting on a vector:
$\mathscr{F}(t,s)\mathscr{a}(s)=\left(\sum_{\ell=0}^{d^{2}-1}\operatorname{V}_{\ell}(t,s)\otimes\overline{\operatorname{V}_{\ell}}(t,s)\right)\mathscr{a}(s)\qquad\forall\,t\,\geq\,s$
The $d^{2}\,\times\,d^{2}$ matrices $\mathscr{F}(t,s)$ are the elements of the semigroup solving the system of differential equations
@@
with generator
@@
$+\sum_{\ell=1}^{\mathrm{d}^{2}-1}\mathscr{c}_{\ell}(t)\left(\operatorname{L}_{\ell}(t)\otimes\overline{\operatorname{L}_{\ell}}(t)-\frac{\operatorname{L}_{\ell}^{\dagger}(t)\operatorname{L}_{\ell}(t)\otimes\mathsf{1}_{d}+\mathsf{1}_{d}\otimes\operatorname{L}_{\ell}^{\top}(t)\overline{\operatorname{L}_{\ell}}(t)}{2}\right)$
where $\top$ denotes matrix transposition. The one-to-one correspondence between (10) and (7) implies that also the operator sum (8) enjoys the semigroup property:
$\Phi(t_{\mathfrak{3}},t_{\mathfrak{2}})\big[\Phi(t_{\mathfrak{2}},t_{\mathfrak{1}})[\operatorname{O}]\big]=\Phi(t_{\mathfrak{3}},t_{\mathfrak{1}})[\operatorname{O}]\qquad\forall\,t_{\mathfrak{3}}\,\geq\,t_{\mathfrak{2}}\,\geq\,t_{\mathfrak{1}}\quad\&\quad\forall\,\operatorname{O}\in\mathcal{M}_{d}(\mathbb{C})$
The semigroup generated by the quantum master equation (5) subject to the conditions (7) always preserves state operators. All these facts combined are reminiscent of the properties of semigroups generated by the classical master equation (1). For this reason quantum master equations satisfying (7) often called “Markovian”. The set of positive maps in quantum theory is, however, far more involved than in the classical case. The positive operator sum form (8) is only a sufficient condition for positivity preservation. In particular, it is equivalent to adding to positivity preservation the requirement [StiW1955, ChoM1975] that the map on $\mathcal{M}_{d^{2}}(\mathbb{C})$ obtained by the tensor product of (8) with the identity on $\mathcal{M}_{d}(\mathbb{C})$ is also positive:
@@
for all positive matrices $\tilde{\operatorname{O}}$ in $\mathcal{M}_{d^{2}}(\mathbb{C})$. The condition (11) is called $d$- or complete positivity. As a consequence, the master equation (1) subject to the Lindblad-Gorini-Kossakowski-Sudarshan conditions (7) may be more descriptively called completely positive rather than Markovian in reference to the property of the solution semigroup [BaNaTh2008].
@@
$\tilde{\Phi}^{(k)}(t,s)[\tilde{\operatorname{O}}]:=(\mathrm{Id}_{k}\otimes\Phi(t,s))[\tilde{\operatorname{O}}]\,\geq\,0\qquad\forall\,t\,\geq\,s$

### v2

@@ Abstract
The spectral properties of positive maps are pivotal for understanding the dynamics of quantum systems interacting with their environment. Furthermore, central problems in quantum information such as the characterization of entanglement can be reformulated in terms of spectral properties of positive maps. The present work aims to contribute to a better understanding of the spectrum of positive maps. Specifically, our main result is a new proof of a universal bound on the $d^{2}-1$ generically non vanishing decay rates $\Gamma_{i}$ of time-autonomous quantum master equations on a $d$-dimensional Hilbert space:
$\phantom{aaaaaaaaaaaaaa}\Gamma_{\mathrm{max}}\,\leq\,\varkappa_{d}\,\sum_{i=1}^{d^{2}-1}\Gamma_{i}$
The prefactor $\varkappa_{d}$ depends only on the dimension $d$ and varies depending on the sub-class of positive maps to which the semigroup solution of the master equation belongs. We provide a brief but self-consistent survey of these concepts. We obtain our main result by resorting to the theory of Lyapunov exponents, a central concept in the study of dynamical systems, control theory, and out-of-equilibrium statistical mechanics. We thus show that progress in understanding positive maps in quantum mechanics may require ideas at the crossroads between different disciplines. For this reason, we adopt a notation and presentation style aimed at reaching readers with diverse backgrounds.
@@ 1 Introduction
Physics is an experimental science. A recurring question is how to extract dynamical laws from experimental data. In the case of open quantum systems, this question is often formulated by asking whether an observed evolution (dynamical map) can be mathematically understood as a family of linear maps that enjoys the semigroup property and whose members are all positivity preserving. In that case, there also exists a precise and computable answer for time-autonomous dynamics [WoEiCuCi2008]. Specifically, one needs to check whether the logarithm of a tomographically estimated dynamical map at a given time, i.e., a single snapshot of the dynamical map, satisfies the necessary and sufficient conditions characterizing generators of completely positive semigroups and, therefore, solutions of the Lindblad [LinG1976] Gorini-Kossakowski-Sudarshan [GoKoSu1976] master equation. These conditions are preservation of the trace and self-adjointness, as well as conditional complete positivity [EvaD1977] (see also [WolM2012, § 7]) of the generator. Whilst the answer is mathematically complete and constructive, actual implementations encounter several stumbling blocks. To start with, the complex logarithm is a multi-valued function so that, in general, checks may require to extend the search to multiple branches. In addition, the search must be robust against state preparation and measurement errors that occur in realistic situations. Finally, the computational cost grows exponentially with the number of dimensions. Connecting to this cost, it is also worth mentioning that it is in fact possible to prove [CuEiWo2012] that the 11-IN-33-SAT, an instance of a NP-complete problem in computational complexity theory [MoMe2011, § 5], is amenable by a Turing reduction to proving that the exponential of a certain non-representational (i.e., abstract and non related to any physical model [IpCa2021]) generator yields an element of a completely positive semigroup. This means that in principle assessing whether the logarithm of a dynamical map specifies the generator of a completely positive semigroup is at least as hard as solving any NP problem on a classical computer. In practice and for systems with few degrees of freedom, the result of [WoEiCuCi2008] is the basis for computationally viable algorithms which are relevant for to-date quantum computing hardware [LiSeKoOnCu2025].
@@
The difficulties discussed above motivate the interest for partial but practically more direct answers that provide necessary conditions (“witnesses”) to identify a completely positive semigroup. It was conjectured long ago [KimG2002] that complete positivity of a homogeneous semigroup of dynamical maps necessarily entails an upper bound on the relaxation rates characterizing the contractive forward evolution. Relaxation rates are commonly measured in laboratory experiments. The conjecture was later extensively corroborated by proofs under additional hypotheses and examples exhibiting tightness of the putative bound [ChKiKoSh2021]. Mathematically, relaxation rates are the real parts of the logarithms of the eigenvalues of the flow solution of a linear differential equation. If the differential equation is time-autonomous, the forward in time flow it generates constitutes a homogeneous semigroup and the relaxation rates coincide with the Lyapunov exponents [AdrL1995]. These considerations led to a first proof of the conjecture by combining ideas of dynamical systems and open quantum system theory [MGKiCh2025]. The proof, however, left open the question of tight bounds for the relaxation rates of positive but non completely positive maps. This is an important question because positive (i.e., positivity-preserving) linear maps—which are fundamental to quantum mechanics—have a far more articulated structure than those appearing in the theory of classical stochastic processes [ChKo2009]. In [ChvEnKiMG2025, vEnChKiMG2025] we were able to provide a complete answer to this question using different and more traditional methods of algebraic quantum theory. Specifically, the results of [ChvEnKiMG2025, vEnChKiMG2025] show that the same bound holds for maps that are at least $2$-positive and that this requirement cannot be weakened further. We emphasize that $2$-positive maps are completely positive maps for single-qubit systems. The relaxation rates of simply positive (or $1$-positive) maps only obey the trivial bound dictated by contractivity, as do classical positive maps. Finally, our results in [ChvEnKiMG2025, vEnChKiMG2025] single out the class of Schwarz maps, a subset of positive maps that contains the class of $2$-positive maps: we show that the relaxation rates of Schwarz maps obey a weaker yet non trivial bound than that of $2$-positive maps.
@@
The structure of the paper is as follows. In section 2 we briefly recall basic facts about positive linear maps in classical and quantum physics. In section 3 we recall that any generator of trace- and self-adjointness preserving linear maps embeds a classical master equation [TeMa1992] if and only if the Kossakowski conditions hold [KosA1972, GoKoSu1976] (see also [RiHu2012, § 4.3]). In section 4 we prove that the foliation of the set of positive maps in quantum mechanics in a hierarchy of nested subsets translates into conditions on the contraction rate of the volume form of state operators. These results are instrumental to prove the main result of this work. In section 5 we briefly review basic concepts of the theory of Lyapunov exponents that we use in the proof of our main result. We give particular emphasis to Lyapunov regularity, a property that quantum master equations relevant for applications, most often time-autonomous or time-(almost)-periodic, enjoy. Section 6 contains our main result. We include some examples illustrating the bounds in section LABEL:sec:examples. The last section is devoted to conclusions and outlook.
@@ 2.1 Classical master equations
Probability vectors valued on a finite-dimensional configuration space are the main objects in the theory of classical stochastic processes. These are real vectors with positive entries that sum to one. Geometrically, they form a simplex, a compact convex subset of a real Euclidean vector space [BeZy2006, § 2].
If the stochastic process is a differentiable Markov chain on a $d$-dimensional Euclidean configuration space, then the family of probability vectors $\mathsf{p}(t)$, that describes the state of the process as $t$ increases, satisfies a first-order system of differential equations of the form
$\dot{\mathsf{p}}(t)=\mathsf{K}(t)\,\mathsf{p}(t)\,.$ (1)
We refer to the family of $d\,\times\,d$-matrices $\mathsf{K}(t)$ as the (time-dependent) generator of the dynamics. We require $\mathsf{K}(t)$ to depend on the time variable $t$ in a sufficiently regular manner to satisfy the conditions of global existence of solutions [AdrL1995, § 1]. In addition, we require $\mathsf{K}(t)$ to be column-diagonally dominated with positive off-diagonal entries, that is,
$\mathsf{K}_{i,j}(t)\begin{cases}\,\geq\,0&\forall\,i\neq j\\[8.5359pt] =-\sum_{k\neq i}^{d}\mathsf{K}_{k,i}(t)&\forall\,i=j\end{cases}$ (2)
@@
$\mathsf{T}(t,s)=\mathcal{T}\exp\left(\int_{s}^{t}\mathrm{d}u\,\mathsf{K}(u)\right)$ (3)
solving (1) with initial condition specified by the identity matrix in $\mathcal{M}_{d}(\mathbb{C})$
$\mathsf{T}(s,s)=\mathsf{1}_{d}$ (4)
@@
$\mathsf{T}(t_{3},t_{1})=\mathsf{T}(t_{3},t_{2})\mathsf{T}(t_{2},t_{1})\qquad\forall\,t_{3}\,\geq\,t_{2}\,\geq\,t_{1}$ (5)
@@
$\|\mathsf{T}(t,s)(\mathsf{p}(s)-\tilde{\mathsf{p}}(s))\|_{1}\,\leq\,\left\|\mathsf{p}(s)-\tilde{\mathsf{p}}(s)\right\|_{1}\,\equiv\,\sum_{i=1}^{d}|\mathsf{p}_{i}(s)-\tilde{\mathsf{p}}_{i}(s)|\,.$ (6)
Combining contractivity with the property of mapping a convex compact set into itself paves the way for the application of Perron-Frobenius theorem (see e.g. [SenE2006, § 1]). The upshot is that the spectrum of any of the elements of the semigroup (3) is confined to the unit disk of the complex plane, and it is symmetric with respect to the real axis. Furthermore, there is at least one eigenvalue equal to one. The right eigenoperator associated with this eigenvalue is a state operator after appropriate normalization.
@@ 2.2 Quantum master equations
State operators are the main objects in the theory of open quantum systems. In the case of a Hilbert space with $d$ dimensions, they are positive semi-definite (henceforth just called “positive”) $d\,\times\,d$ matrices with unit trace. Geometrically, they also form a compact and convex subspace of the matrix space $\mathcal{M}_{d}(\mathbb{C})$ of $d\,\times\,d$ matrices with complex entries [BeZy2006, § 8.4].
Based only on the requirements of trace and self-adjointness preservation, it is in general possible to derive the matrix valued equation
$\dot{\bm{\rho}}(t)=\mathfrak{L}(t)[\bm{\rho}(t)]\,.$ (7)
@@
$\begin{split}&\mathfrak{L}(t)[\operatorname{O}]=-\imath\,\left[\operatorname{H}(t),\operatorname{O}\right]+\frac{1}{2}\sum_{\ell=1}^{d^{2}-1}\mathscr{c}_{\ell}(t)\Big(\left[\operatorname{L}_{\ell}(t)\,,\operatorname{O}\operatorname{L}_{\ell}^{\dagger}(t)\right]+\left[\operatorname{L}_{\ell}(t)\operatorname{O}\,,\operatorname{L}_{\ell}^{\dagger}(t)\right]\Big)\\ &\begin{cases}\operatorname{Tr}\left(\operatorname{L}_{\ell}^{\dagger}(t)\operatorname{L}_{k}(t)\right)=\delta_{\ell,k}\\ \operatorname{Tr}\left(\operatorname{L}_{\ell}(t)\right)=0\\ \operatorname{H}(t)=\operatorname{H}^{\dagger}(t)\end{cases}\hskip 14.22636pt\forall\,t\,\geq\,0\hskip 14.22636pt\&\hskip 14.22636pt\forall\,\ell,k=1,\dots,d^{2}-1\end{split}$ (8)
where $\operatorname{H}(t)$ is a Hamilton operator, the $\mathscr{c}_{\ell}(t)$’s are scalar couplings, and the $\operatorname{L}_{\ell}(t)$’s are commonly referred to as decoherence or jump operators. As in the classical case, here we always assume that any explicit time dependence is sufficiently regular, e.g. continuous and uniformly bounded, to satisfy the conditions for global existence of solutions of (7) for any positive time $t\,\geq\,0$.
We refer to (7), always implying the generator form (8), as “quantum master equation”. The main problem with (7) is that it does not ensure the preservation of positivity without further assumptions on the scalar couplings $\mathscr{c}_{\ell}(t)$.
A cornerstone result of the theory of open quantum systems is the proof, see e.g. [RiHu2012, Theorem 4.2.1], that if and only if
$\mathscr{c}_{\ell}(t)\,\geq\,0\qquad\forall\,\ell=1,\dots,d^{2}-1$ (9)
then solutions of (7) can always be couched in the readily positivity-preserving operator sum form (often referred to as the Kraus form [KraK1983])
$\Phi(t,s)[\bm{\rho}(s)]=\sum_{\ell=0}^{d^{2}-1}\operatorname{V}_{\ell}(t,s)\bm{\rho}(s)\operatorname{V}_{\ell}^{\dagger}(t,s)\qquad\forall\,t\,\geq\,s$ (10)
@@
$\sum_{\ell=0}^{d^{2}-1}\operatorname{V}_{\ell}^{\dagger}(t,s)\operatorname{V}_{\ell}(t,s)=\operatorname{1}_{d}\qquad\forall\,t\,\geq\,s$ (11)
@@
The proof of (9)-(10) was first obtained in groundbreaking papers by Lindblad [LinG1976] and Gorini-Kossakowski-Sudarshan [GoKoSu1976] under the hypothesis of a time-autonomous generator. This means the case where in (8) the Hamilton operator $\operatorname{H}$, the scalar couplings $\mathscr{c}_{\ell}$’s, and decoherence operators $\operatorname{L}_{\ell}$’s are all time independent. In the time-autonomous case (9) straightforwardly reduces to the requirement that the $\mathscr{c}_{\ell}$’s are positive real numbers:
$\dot{\bm{\rho}}(t)=\mathfrak{L}[\bm{\rho}(t)]:=-\imath\,\left[\operatorname{H},\bm{\rho}(t)\right]+\frac{1}{2}\sum_{\ell=1}^{d^{2}-1}\mathscr{c}_{\ell}\Big(\left[\operatorname{L}_{\ell}\,,\bm{\rho}(t)\operatorname{L}_{\ell}^{\dagger}\right]+\left[\operatorname{L}_{\ell}\bm{\rho}(t)\,,\operatorname{L}_{\ell}^{\dagger}\right]\Big),\ \ c_{\ell}\geq 0.$ (12)
As a consequence, solutions of the master equation only depend upon the time elapsed from the moment an initial state operator $\bm{\rho}$ is assigned:
$\bm{\rho}(t)=\Phi(t,0)[\bm{\rho}],$ (13)
and
$\Phi(t,s)=\Phi(t-s,0)=e^{(t-s)\mathfrak{L}}.$ (14)
Correspondingly, we can regard the collection of linear maps $\Phi$ and matrices $\operatorname{V}_{\ell}$’s as forming a one parameter family, also called homogeneous, with respect to the time variable. The time-autonomous case is of particular interest for applications because it describes quantum master equations obtained in the van-Hove scaling limit [DavE1974].
It is worth recalling that we can always couch the matrix equation (7) in the form of a vector equation using the vectorization isomorphism which relates the outer and tensor products [BeZy2006, § 8.2]. Accordingly, a solution of (7) is in a one-to-one correspondence with a family of vectors in $\mathbb{C}^{d^{2}}=\mathbb{C}^{d}\otimes\mathbb{C}^{d}$
$\bm{\rho}(t)=\sum_{i,j=1}^{d}\mathbf{e}_{i}^{\scriptscriptstyle{(d)}}\mathbf{e}_{j}^{\scriptscriptstyle{(d)}}{}^{\dagger}\rho_{i,j}(t)\quad\Longleftrightarrow\quad\mathscr{r}(t)=\sum_{i,j=1}^{d}\mathbf{e}_{i}^{\scriptscriptstyle{(d)}}\otimes\mathbf{e}_{j}^{\scriptscriptstyle{(d)}}\rho_{i,j}(t)$ (15)
where $\left\{\mathbf{e}_{i}^{\scriptscriptstyle{(d)}}\right\}_{i=1}^{d}$ is the canonical (or computational) basis of $\mathbb{C}^{d}$. In other words, (15) says that we can construct a vector from a matrix by stacking rows one above the other. Upon applying this prescription and using the overline symbol $\overline{(\cdot)}\,$ to denote entry-wise complex conjugation, the evolution map (10) takes the familiar form in linear dynamical systems theory (see, e.g., [CeCeVu2009]) of a matrix acting on a vector:
$\mathscr{F}(t,s)\mathscr{r}(s)=\left(\sum_{\ell=0}^{d^{2}-1}\operatorname{V}_{\ell}(t,s)\otimes\overline{\operatorname{V}_{\ell}}(t,s)\right)\mathscr{r}(s)\qquad\forall\,t\,\geq\,s$ (16)
For any fixed values of $t$ and $s$, $\mathscr{F}(t,s)$ is a $d^{2}\,\times\,d^{2}$ complex square matrix:
$\mathscr{F}(t,s)\in\mathcal{M}_{d^{2}}(\mathbb{C})=\mathcal{M}_{d}(\mathbb{C})\otimes\mathcal{M}_{d}(\mathbb{C})$ (17)
As we vary $t$ and $s$, the $\mathscr{F}(t,s)$’s specify the elements of the semigroup solving the “reshaped” master equation [BeZy2006, § 10.2]
@@
Here, $\mathsf{1}_{d^{2}}$ is the identity matrix in $\mathcal{M}_{d^{2}}(\mathbb{C})$ and
@@
$+\sum_{\ell=1}^{d^{2}-1}\mathscr{c}_{\ell}(t)\left(\operatorname{L}_{\ell}(t)\otimes\overline{\operatorname{L}_{\ell}}(t)-\frac{\operatorname{L}_{\ell}^{\dagger}(t)\operatorname{L}_{\ell}(t)\otimes\mathsf{1}_{d}+\mathsf{1}_{d}\otimes\operatorname{L}_{\ell}^{\top}(t)\overline{\operatorname{L}_{\ell}}(t)}{2}\right)$ (19)
We use the superscript $\top$ to denote matrix transposition. The one-to-one correspondence between (18) and (9) implies that the operator sum (10) also enjoys the semigroup property:
$\Phi(t_{3},t_{2})\big[\Phi(t_{2},t_{1})[\operatorname{O}]\big]=\Phi(t_{3},t_{1})[\operatorname{O}]\qquad\forall\,t_{3}\,\geq\,t_{2}\,\geq\,t_{1}\quad\&\quad\forall\,\operatorname{O}\in\mathcal{M}_{d}(\mathbb{C})$ (20)
The semigroup generated by the quantum master equation (7) subject to the conditions (9) always preserves the state operators. All these facts combined are reminiscent of the properties of semigroups generated by the classical master equation (1). For this reason quantum master equations satisfying (9) are often called “Markovian”. The set of positive maps in quantum theory is, however, much more involved than in the classical case. The positive operator sum form (10) is only a sufficient condition for positivity preservation. In particular, (10) is equivalent to adding to positivity preservation the requirement [StiW1955, ChoM1975] that the map on $\mathcal{M}_{d^{2}}(\mathbb{C})$ obtained by the tensor product of (10) with the identity on $\mathcal{M}_{d}(\mathbb{C})$ is also positive:
@@
for all positive matrices $\tilde{\operatorname{O}}$ in $\mathcal{M}_{d^{2}}(\mathbb{C})$. The condition (21) is called $d$- or complete positivity. As a consequence, the master equation (7) subject to the Lindblad-Gorini-Kossakowski-Sudarshan conditions (9) may be more descriptively called completely positive rather than Markovian with reference to the property of the solution semigroup [BaNaTh2008].
@@
$\tilde{\Phi}^{(k)}(t,s)[\tilde{\operatorname{O}}]:=(\mathrm{Id}_{k}\otimes\Phi(t,s))[\tilde{\operatorname{O}}]\,\geq\,0\qquad\forall\,t\,\geq\,s$ (22)